# TAViS: Text-bridged Audio-Visual Segmentation with Foundation Models

### v1

@@ 4.3 Ablation Study
To further demonstrate the effectiveness of using text as a bridge, we employ t-SNE [39] to visualize the audio class and visual class embeddings under different settings. For clearer visualization, we focus on the S4 dataset results, which contains single-class images. We specifically exclude classes with insufficient samples, as it hard to cluster. As illustrated in Figure 4, utilizing text as an intermediate bridge leads to more compact clusters within each class, effectively mitigating the impact of modality-specific noise. In contrast, when removing either text-related loss term ($\mathcal{L}_{a2t}$ or $\mathcal{L}_{i2t}$), the direct audio-visual alignment results in more dispersed clusters, indicating the challenge of handling noisy representations without textual guidance.
@@
Figure 5: Qualitative comparison of our model against state-of-the-art AVSS methods. From top to bottom: (1) audio visibility graphs; (2) visual frames; (3) prediction maps from AVSegFormer [9]; (4) prediction maps from SAMA-AVS [27] (left) and AVSBench [45] (right); (5) prediction maps from our model; and (6) ground truth.
@@ 4.4 Comparison with State-of-the-Art Methods
We report the performance comparison of our methods against top-performing state-of-the-art methods, including 14 AVS methods [45, 13, 21, 29, 30, 25, 9, 26, 22, 44, 18, 33, 27, 40]. As shown in Table 3, we first compare performance on the S4 and MS3 datasets, and our model demonstrates consistent improvements over existing methods on both datasets. Furthermore, we evaluate our model on the AVSS dataset by computing the similarity between audio embedding $T_{a}$ and text embedding $T^{t}$ for class prediction of segmented regions. Our approach not only surpasses existing methods on AVSS but also extends SAM-based methods with class prediction capabilities. Notably, increasing the image size from 224 to 1024 significantly improves performance, highlighting the importance of image resolution in SAM(SAM2)-based models, as 1024 matches the foundation model’s original training resolution. Figure 5 provides visual comparisons among the top-performing models.
@@ 4.5 Analysis of Zero-shot Generalization Ability
Previous AVSS methods typically predict classes by applying an MLP operation on a fixed set of classes [44, 32], which limits their ability to generalize in real-world scenarios. In contrast, our model leverages ImageBind, which was trained on a broader set of datasets that extend beyond single audio-visual datasets. This enables superior generalization to unseen classes. Following the approach and setting in [11], we train the model on seen classes and evaluate it on unseen classes. To predict the specific class, we compute the similarity of text embeddings with audio and image embeddings respectively, combine them with an addition operation (coefficient set to 1) to obtain the final class prediction. As shown in Table 6, our model outperforms existing OV-AVSS methods with less trainable parameters on zero-shot setting, highlighting the effectiveness of aligning foundation models. While more specialized AVS foundation models may be better suited for the AVSS task [32], the lack of text alignment in such models hinders their generalization ability.

### v2

@@ 4.3 Ablation Study
To further demonstrate the effectiveness of using text as a bridge, we employ t-SNE [van2008visualizing] to visualize the audio class and visual class embeddings under different settings. For clearer visualization, we focus on the S4 dataset results, which contains single-class images. We specifically exclude classes with insufficient samples, as it hard to cluster. As illustrated in Figure 4, utilizing text as an intermediate bridge leads to more compact clusters within each class, effectively mitigating the impact of modality-specific noise. In contrast, when removing either text-related loss term ($\mathcal{L}_{a2t}$ or $\mathcal{L}_{i2t}$), the direct audio-visual alignment results in more dispersed clusters, indicating the challenge of handling noisy representations without textual guidance.
@@
Figure 5: Qualitative comparison of our model against state-of-the-art AVSS methods. From top to bottom: (1) audio visibility graphs; (2) visual frames; (3) prediction maps from AVSegFormer [gao2024avsegformer]; (4) prediction maps from SAMA-AVS [liu2024annotation] (left) and AVSBench [zhou2023audio] (right); (5) prediction maps from our model; and (6) ground truth.
@@ 4.4 Comparison with State-of-the-Art Methods
We report the performance comparison of our methods against top-performing state-of-the-art methods, including 14 AVS methods [zhou2023audio, hao2024improving, li2023catr, mao2023contrastive, mao2023multimodal, liu2024bavs, gao2024avsegformer, liu2024audio, li2024qdformer, yang2024cooperation, kirillov2023segment, mo2023av, liu2024annotation, wang2024prompting]. As shown in Table 3, we first compare performance on the S4 and MS3 datasets, and our model demonstrates consistent improvements over existing methods on both datasets. Furthermore, we evaluate our model on the AVSS dataset by computing the similarity between audio embedding $T_{a}$ and text embedding $T^{t}$ for class prediction of segmented regions. Our approach not only surpasses existing methods on AVSS but also extends SAM-based methods with class prediction capabilities. Notably, increasing the image size from 224 to 1024 significantly improves performance, highlighting the importance of image resolution in SAM(SAM2)-based models, as 1024 matches the foundation model’s original training resolution. Figure 5 provides visual comparisons among the top-performing models.
@@ 4.5 Analysis of Zero-shot Generalization Ability
Previous AVSS methods typically predict classes by applying an MLP operation on a fixed set of classes [yang2024cooperation, mo2024unveiling], which limits their ability to generalize in real-world scenarios. In contrast, our model leverages ImageBind, which was trained on a broader set of datasets that extend beyond single audio-visual datasets. This enables superior generalization to unseen classes. Following the approach and setting in [guo2024open], we train the model on seen classes and evaluate it on unseen classes. To predict the specific class, we compute the similarity of text embeddings with audio and image embeddings respectively, combine them with an addition operation (coefficient set to 1) to obtain the final class prediction. As shown in Table 6, our model outperforms existing OV-AVSS methods with less trainable parameters on zero-shot setting, highlighting the effectiveness of aligning foundation models. While more specialized AVS foundation models may be better suited for the AVSS task [mo2024unveiling], the lack of text alignment in such models hinders their generalization ability.